Q:- A variable plane passes through a fixed point (a,b,c) and meets the coordinate axes in A,B,C . Locus of the point common to the planes through A,B,C and parallel to coordinate plane is...............


 


A. a/x+b/y+c/z=1


B.x/a+y/b+z/c=1


C.ax+by+cz=1


D. None of these

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

let equation of plane is

x/a1 + y/b1 + z/c1 =1

then co ordinate of A is (a1,0,0)  ,B is (0,b1,0) ,C is (0,0,c1)


this plane passes through (a,b,c)

so

 a/a1 +b/b1 + c/c1  =1 .....................(1)

equation of plane passes trough A and parallel to coordinate plane  is

 x= a1

equation of plane passes trough B and parallel to coordinate plane  is

 y= b1

equation of plane passes trough C and parallel to coordinate plane  is

 z= c1

let common poin in these plane is (x1,y1,z1)

 so x1=a1

     y1=b1

     z1=c1

put value of a1,b1,c1 in equation 1

a/x1 + b/y1 + c/z1 =1


so locus is

 a/x +b/y + c/z =1


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin


Sachin Tekale
5 years ago
A variable plane is at a constant distance 3p from the origin and meets the axes in A,B and C . Show that the locus of the centroid of the triangle ABC is x-2+y-2+z-2=p-2